INAF queue observing, 2015 Feb 13-14

Observer: F. Cusano, R. Carini
Telescope Operator: S. Allanson
Support Astronomer: Olga Kuhn on call
Instrument: LUCI/MODS/LBC

Summary

At the sunset the sky is clear

Data Summary

Project_RUN Objects Filter ExpTime Notes
Calibration MODS for Israel program
telluric HIP 41798 HK    
Rossi CD GRB020127 HK 44 m  
Mignoli MOS 20 J1411+12 red grating 2 h  
Mignoli LBC MOS20_red J1411+12 filter z-SLOAN 15 m  
Mignoli LBC MOS20_red J1411+12 filter i-SLOAN 30 m  
Mignoli LBC MOS20_red J1411+12 filter r-SLOAN 30 m  
standard Hz44 red grating    

Details

12.17 flat calibration for Israel Program mods1b.20150214.0001-0005, mods1r.20150214.0001-0005

12:30 arc calibration for Israel Program mods1b.20150214.0006, mods1r.20150214.0006
  • in the blue channel the spectra is saturated, we modify the script with time exposure 1 s mods1b.20150214.0007, mods1r.20150214.0007

01:00 TO opens the dome

02:03 star the acquisition of telluric star for Rossi CD program executeLUCIScript.sh GRB020127telluric_ls.acq luci.20150214.0006

02.12 executeLUCIScript.sh GRB020127telluric_ls.spec luci.20150214.0007-0011
  • seeing 1.5

02.25 slew to the target of Rossi CD executeLUCIScript.sh GRB020127_ls.acq luci.20150214.0012-0016

02:37 executeLUCIScript.sh GRB020127_ls.spec luci.20150214.0017-00
  • 2:59 seeing 0.6
  • 3:18 seeing 1.1
  • 3:22 cloud coming in, images -0024, 0025, 0026 maybe not good for clouds
  • 3:31 guide star fainter of 1 mag due to clouds, seeing 0.9

3:40 abort script due to clouds.

10:17 the conditions are better. we slew to Mignoli MOS20 program

10.27 acqMODS mods.20..acq mods1r.20150214.0008-0010
  • dofpia, /first, /X2, /redonly

10:30 run 4_MOS20_z.ob

10:40 execMODS mods.20.3.obs mods1r.20150214.0011-0014
  • seeing 0.9"

10:49 run 5_MOS20_i.ob, lbcr.20150214.105726.fits done during the offset of MODS
  • seeing 0.8"

11:28 dofpia, /X2, /redonly

11:31 we get one more image in i-SLOAN filter to recover the lost one.
  • run 5_MOS20_i.ob

11:38 run 6_MOS20_r.ob

11.51 10:40 execMODS mods.20.4.obs mods1r.20150214.0015-0018
  • seeing 0.8"

12:22 dofpia, /X2, /redonly

12:32 run 7_MOS20_z.ob
  • LBC red channel get stuck, one one image has been done

1:02 slew to standard Hz44 acqMODS hz44.acq mods1r.20150214.0019
  • expecMODS hz44.obs mods1r.20150214.0020-22
  • standard in twilight, seeing 1.1"

-- %USERSIG{RobertaCarini - 2015-02-14}%

Comments

 
Topic revision: r7 - 19 May 2015, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback